OpenStudy (brightemmerald):

What number must you add to complete the square? x^2 + 12x = -5 A. 144 B. 36 C. 6 D. 12

OpenStudy (jackthegreatest):

did you see what I wrote earlier?

OpenStudy (brightemmerald):

i accidentally deleted it

OpenStudy (studygurl14):

ax^2 + bx + c (b/2)^2 = the number you add to complete the square

OpenStudy (precal):

take 12/2 and then square it

OpenStudy (jackthegreatest):

or well the only square that will get you 12x is (x+6)(x+6)

OpenStudy (mathstudent55):

When the coefficient of the x^2 term is 1, add the square of half of the coefficient of x.

OpenStudy (brightemmerald):

That was a lot of similar information at once.

OpenStudy (jdaub18):

B, 36.

OpenStudy (3mar):

\[x^2+12x=-5\] we need this form to be in that form: \[x^2+12x+(\frac{ 12 }{ 2 })^2-(\frac{ 12 }{ 2 })^2=-5\] then complete the square \[[x^2+12x+36]-36=-5\] \[(x+6)^2-36+5=0\] \[(x+6)^2-31=0\]

OpenStudy (triciaal):

are those the correct options?

OpenStudy (brightemmerald):

Are you positive? jdaub18

OpenStudy (mathstudent55):